OpenStudy (anonymous):

If f(x) varies directly with x2, and f(x) = 96 when x = 4, find the value of f(2). 192 72 32 24

OpenStudy (math&ing001):

f(x) varies directly with x2 means f(x) = a*x2. You also got f(4)=a*4^2=16a=96 So a=6 therefore f(2) = a*2^2 = 6*4 = 24

OpenStudy (anonymous):

thanks :)

OpenStudy (math&ing001):

welcome !